# Link Me Baby One More Time: Social Music Discovery on Spotify

### v1

@@ Abstract
We explore the social and contextual factors that influence the outcome of person-to-person music recommendations and discovery. Specifically, we use data from Spotify to investigate how a link sent from one user to another results in the receiver engaging with the music of the shared artist. We consider several factors that may influence this process, such as the strength of the sender-receiver relationship, the user’s role in the Spotify social network, their music social cohesion, and how similar the new artist is to the receiver’s taste. We find that the receiver of a link is more likely to engage with a new artist when (1) they have similar music taste to the sender and the shared track is a good fit for their taste, (2) they have a stronger and more intimate tie with the sender, and (3) the shared artist is popular with the receiver’s connections. Finally, we use these findings to build a Random Forest classifier to predict whether a shared music track will result in the receiver’s engagement with the shared artist. This model elucidates which type of social and contextual features are most predictive, although peak performance is achieved when a diverse set of features are included. These findings provide new insights into the multifaceted mechanisms underpinning the interplay between music discovery and social processes.
@@ Introduction
Spotify has over 574 million users, and over 100 million available tracks (Spotify 2023). The platform also offers a social aspect; users have various ways of interacting with others on Spotify, from following the activity of their friends, sharing music, to collaboratively building playlists together.
@@
The Spotify social network comprises its users (nodes) and their interactions on the platform (edges). Each edge has an associated timestamp, indicating when an interaction occurred. This network is multiplex; the edges represent three different types of interactions; collaborative playlists, social listening sessions and link shares. A collaborative playlist is a special type of playlist that can be modified by a group of users. Although these are group interactions, we represent them as multiple pairwise interactions (edges) between each pair of collaborating users. The edges representing these interaction are un-directed, as they are a reciprocal interaction. A social listening session allows multiple users to add songs to a real-time queue, listening to the same music either remotely or together. Like collaborative playlists, these are reciprocal interactions, and the edges are un-directed. A link share is a link to a track, album or artist sent from one user to another via a messaging or social media platform. This interaction is directed, and is represented in the network by an edge from the sender to the receiver. We construct the social network by aggregating the relevant social interactions (collaborative playlists, social listening sessions and link shares) occurring on the platform prior to the study period, weighting each interaction type equally. The result is a weighted, directed network with $\sim$2.2B edges between $\sim$518M users (nodes).
@@
We only consider “discovery” share events where the receiver has not listened to the artist before and we filter the data to include only link shares of tracks for which 1) the link was opened (comprising 50% of all shares), and 2) the receiver’s Spotify account registered a playback event of the shared content for at least 30 seconds (64% of all opened shares). This ensures that the receiver did indeed come into contact with the music that was shared. To account for popularity bias, we randomly sample up to 1M share events stratified by artist across various popularity bins. Our final sample consists of $\sim$2.72M events across 3,102 distinct artists between April-June 2023. Each event contains anonymised sender and receiver identifiers, the shared track metadata as well as artist popularity and album age at the time of the share. In addition, we have metadata such as the relevant timestamps and the application to which the link was shared. Each share of a track link is associated with the track’s primary artist in the case of multiple artists.
We split the social media application type into two categories, direct or broadcast, depending on whether they represent one-to-one or one-to-many modes of communication. The applications we identify as direct are: WhatsApp, Facebook Messenger, SMS, Line, Instagram Direct, and Samsung Messenger. The applications we identify as broadcast are: Instagram Stories, Facebook Feed, Facebook Stories, and X, formerly Twitter. Some sharing events have an unknown destination platform or fit into neither category. These events are excluded from the analysis where share category is relevant. Note that the applications we identify as one-to-one can also be used to message group chats, but the key distinction here is in the magnitude of number of people receiving the link.
@@
To this end, we use vector representations of Spotify users and content in latent space. These vector embeddings are computed daily on Spotify by training the word2vec algorithm (Mikolov et al. 2013) on user generated playlists, with the playlists as the “documents” and the tracks as the “words”. This process results in 80-dimensional embedding vectors for each track, where tracks that are often co-listed in playlists appear close together in the embedding space. Users are also embedded in this space by taking the average of vectors over all tracks they have listened to over a certain time window. This embedding space is consistent with the way that users engage with music (Mok et al. 2022; Anderson et al. 2020). We use vectors computed at the beginning of our analysis period for each user and track entity in our sample, to ensure consistency. The vectors can be used to understand the taste similarity of two users, or between a user and track. We define this as the cosine similarity between the two vectors.
@@
To define the success of a share, we are interested in a user’s engagement with an artist before and after a share event. We denote the number of tracks that a user $i$ listens to for longer than 30s by artist $\alpha$ on a particular day $t$ as $n_{i,\alpha}(t)$. We then define the user-artist engagement as $e_{i,\alpha}(t)=\log_{10}(n_{i,\alpha}(t))+1$, if $n_{i,\alpha}(t)\geq 1$, and 0 otherwise.
User-artist engagement captures the intensity of the user’s engagement with the artist each day, which we aggregate over time as:
@@
where $t_{0}$ is the day when the link share is opened, and $K$ is the number of days over which we aggregate, which can be either positive or negative indicating aggregation over time before or after $t_{0}$ respectively. We use this metric to define the past engagement of a sender and the future engagement of a receiver with the shared artist. We define the success of a share event as the binary outcome of the receiver’s future engagement with the shared artist.
To this end, we define an engaged receiver as anyone who has listened to two or more tracks by the artist in the 7 days following the share, this is equivalent to a threshold of $E_{i,\alpha}(t_{0},7)>1.3$, since $\log_{10}{2+1}=1.3$. This choice of threshold identifies receiving users who are beginning to form an attachment with the artist in the short-term and are becoming potential fans. By this definition, $47.2\%$ of link shares result in successful engagement. Note that there are other choices that could be made to define a successful share. In particular, here we are interested in behaviour change on a short-term timescale directly following a share.
@@ Results
We evaluate the three hypotheses set out at the beginning of the paper by examining the probability of engagement $p(engaged)$, i.e. the proportion of link share events where the receiver’s engagement is $E_{i,\alpha}(t_{0},7)>1.3$ out of all share events, with respect to: the music taste similarity of the pair of sharing users and the shared track (H1), their relationship strength (H2), the social cohesion of their neighbourhood with respect to an artist (H3). We then construct and evaluate a predictive model of new listener engagement, using a broader set of social and contextual features.
@@
We measure music taste similarity as the cosine similarity of the music vectors described in Data and Methods. We first observe the level of homophily present in the sample of people sharing links, whether successful or not. We randomly shuffle the sender in user pairs in our sample, creating random permutation pairs of the social network, and take the user pair cosine similarity of their taste vectors. Figure 1 compares the distribution of the user pair cosine similarities between random pairs, and between the pairs who have shared and opened links, where we note that the distribution of user pair cosine similarities for users involved in link sharing is skewed towards higher similarity with a mean of 0.54, compared to the random pairs with a mean of 0.11. This gives some indication of the level of homophily in the Spotify social network; either users are more likely to become friends with other users who share similar tastes, or they are more likely to share with their friends who have similar tastes.
For this reason, it is difficult to determine whether the engagement we observe following a share is due to homophily or social contagion. However, while we may not be able to distinguish between the two mechanisms at play, we can still investigate the social and contextual factors that lead to engagement as a new listener of the shared artist. We first test the hypothesis (H1) that the probability of engagement will increase when users have similar taste in music (sender-receiver cosine similarity), or when the shared track is similar to the listening history of the receiver (receiver-track cosine similarity). The results are given in Figure 2 (a), showing the distribution of sender-receiver cosine similarity in our sample, and $p(engaged)$ as a function of the similarity. Similarly, Figure 2 (b) shows the results for receiver-track cosine similarities. We observe that while (H1) is confirmed, there is a stronger effect for the receiver-track cosine similarity; $p(engaged)$ increases linearly with cosine similarity.
@@
We also investigate the effect of the sender-artist engagement for the artist on $p(engaged)$. We measure this by taking the 7 day aggregate user-artist engagement (described in Data and Methods) for the week prior to the share, which is a signal for how engaged the sender was with the artist, prior to sharing their music. This corresponds to Equation 1, where $i$ is the sender, $t_{0}$ is the day of the share, and $K=-7$. This feature can be thought of as a proxy for the strength of the sender’s recommendation. Users have the ability to add messages when sharing a Spotify link, and while we do not have access to this textual data, it is possible that the associated message might be more emphatic when the sender is more engaged with the artist. These results are shown in Figure 3. The probability of engagement of the receiver increases with the sender-artist engagement.
@@
As suggested in (Almaatouq et al. 2016) and (Bond et al. 2012), stronger social ties are more effective at spreading behaviours. We explore the hypothesis (H2) that the probability of engagement $p(engaged)$ will increase with stronger and more intimate ties, as represented by the tie strength from the social network. We measure tie strength in three ways; reciprocity, the mode of social media application used to share the track, and sum of social interactions. Reciprocity, is a boolean variable that indicates whether the receiver has ever shared a link with the sender, prior to share time. Figure 4 (a) shows the distribution of sender-receiver cosine similarities over all share events, separated by reciprocal and non-reciprocal ties. For reciprocal links, the mean sender-receiver cosine similarity is 0.58, compared to 0.5 for non-reciprocal links, indicating that users connected by reciprocal ties tend to have a greater taste similarity than those who do not. Figure 4 (b) shows the probability of becoming an engaged receiver, $p(engaged)$, along a reciprocal and non-reciprocal link, holding constant the receiver-track cosine. Even accounting for the taste of the receiver, the probability of future engagement is more likely along a reciprocal tie. There is noise at very large and very small receiver-track cosine similarities, where there is little data available.
@@
Figure 4 (c) shows the distribution of sender-receiver cosine similarities over all share events, separated by ties that represent direct or broadcast platform share events. For the users sharing via direct social media applications, the mean user-user cosine similarity is 0.54, compared to the 0.57 for users sharing via broadcast methods. This indicates that users selectively engage with music shared by friends with similar taste on social media. Figure 4 (d) shows the probability of becoming engaged with the artist for share events via direct and broadcast mediums, again holding constant the user-track cosine similarity. Accounting for the taste of the receiver, the future engagement with an artist is more likely to occur when the mode of communication is a direct one, even when the track is not as good a match for the receiver’s taste. Note that the probability of engagement goes almost to 0 for broadcast application types at large receiver-track cosine similarities.
@@
The final measure of tie strength, sum of social interactions, the number of times that the sender and receiver have interacted before the share on Spotify. This includes link shares (from the sender to receiver or from the receiver to the sender), social listening sessions and blends. This measure gives a sense of whether the social tie is new, or previously established. Figure 5 shows the distribution of the sum of social interactions over the pairs of users in our sample, as well as the probability of engagement along a link as a function of the sum of social interactions. The probability of of an engagement jumping from 0.44 to 0.48 just after one previous interaction and then gradually up to 0.5, until 5 interactions. After 5 interactions, the probability of engagement remains stable. This indicates that while engagement is more likely to happen along an established tie, there is a “saturation point” at which more previous interaction does not increase the probability.
@@
Next, we explore the effects of social cohesion. These results investigate how the level of engagement with the shared artist in the receiver’s neighbourhood of contacts affects the probability of becoming engaged as a receiver of a share link (H3). We define the receiver’s “friends” as anyone with whom they have a reciprocal relationship in the social network. Figure 6 (a) shows the distribution of number of friends over receivers in the sample. The average number of friends is 3.95, but the distribution is skewed to the left, most users have 0 or 1 friend. A receiver’s friend is considered to be engaged with the shared artist if they have a user-artist engagement indicative of listening to at least one track per day on average by the shared artist in the six months prior to the analysis period. This is equivalent to the condition $E_{i,\alpha}(t_{0},180)>180$, where $i$ is the friend user, $\alpha$ is the shared artist and $t_{0}$ is the start of the analysis period. We expect that the probability of engagement of a receiving user will increase with the number of friends that the receiver has, who are already engaged with the artist, as a consequence of either social influence, or homophily in the social network.
We group all receivers who have the same number of friends, and track the probability of engagement as a function of the number of engaged friends of the receiver. The reason for holding constant the number of friends is to control for the absolute number of friends, as well as the fraction who are engaged. The results are shown in Figure 6 (b) for 2,3 and 5 friends. The shaded area gives the $95\%$ confidence interval. In all cases, the probability of engagement increases with the number of engaged friends, supporting the hypothesis that social cohesion increases the probability of engagement.
It is possible that this effect that we observe is related to the artist popularity. In particular, the fraction of listener friends who are engaged with an artist may simply be a proxy for popularity, as these artists have more engagement overall. To disentangle these effects, we separate out the artist according to their popularity on Spotify, and plot the probability of engagement as a function of number of engaged friends. The results are shown in Figure 6 (c). The probability of engagement increases with the number of engaged friends, even when accounting for artist popularity. More interestingly, we notice that this effect is more pronounced for the less popular artists, perhaps indicating that such social effects play a more important role in music discovery for lesser known artists.
@@
We would also like to evaluate the insights from H1-3 together with other contextual factors which might influence short-term engagement following a share. To this end, we construct a Random Forest classifier to help us distinguish between factors of varying strength for the prediction of whether a shared track will result in engagement with the shared artist in the week following the share. This model elucidates further the social and contextual factors that are important for a successful person-to-person music recommendation.
@@
Sum of social interactions
Sum of the number of social interactions between the sender and receiver prior to share-time. Includes collaborative playlist additions, social listening sessions, and link shares between the two users in either direction.
Direct link share
Boolean, True if the social media application type used to share the link was direct (e.g. Whatsapp), as opposed to broadcast or unknown.
Reciprocal link sharing
Boolean, True if receiver has ever shared a link with the sender prior to share.
@@
Receiver in-degree
Number of shares received by the receiver (from anyone) prior to share-time.
Receiver out-degree
Number of shares sent by receiver prior to share-time.
Sender out-degree
Number of shares sent by the sender before the share-time.
@@
Fraction of engaged friends
“Friend” is defined as someone with whom the receiver has a reciprocal relationship on social network prior to share-time. A friend is considered engaged if $E_{i,\alpha}(t_{0},180)>180$, where $i$ is the friend user, $\alpha$ is the shared artist and $t_{0}$ is the start of the analysis period.
@@
sender-receiver cosine similarity
Cosine similarity between receiver and sender.
receiver-track cosine similarity
Cosine similarity between receiver and track.
@@
Artist popularity at share-time
Rank of the primary artist in Spotify by number of streams at the time of share.
Release age at share-time
Time elapsed between album release time and share-time (seconds).
@@
Sender-artist engagement 7 days prior
Sender-artist engagement over 7 days prior to share-time.
@@
is subscriber
Whether the user was a subscriber when the share was opened.
receiver streaming 7 days prior
Number of hours of streamed music in the previous 7 days to opening the share.
receiver day on platform
Number of days passed since the user registered to the platform to the day the share was opened.

### v2

@@ Abstract
We explore the social and contextual factors that influence the outcome of person-to-person music recommendations and discovery. Specifically, we use data from Spotify to investigate how a link sent from one user to another results in the receiver engaging with the music of the shared artist. We consider several factors that may influence this process, such as the strength of the sender-receiver relationship, the user’s role in the Spotify social network, their music social cohesion, and how similar the new artist is to the receiver’s taste. We find that the receiver of a link is more likely to engage with a new artist when (1) they have similar music taste to the sender and the shared track is a good fit for their taste, (2) they have a stronger and more intimate tie with the sender, and (3) the shared artist is popular amongst the receiver’s connections. Finally, we use these findings to build a Random Forest classifier to predict whether a shared music track will result in the receiver’s engagement with the shared artist. This model elucidates which type of social and contextual features are most predictive, although peak performance is achieved when a diverse set of features are included. These findings provide new insights into the multifaceted mechanisms underpinning the interplay between music discovery and social processes.
@@ Introduction
Spotify has over 574 million users, and over 100 million available tracks (Spotify 2023). The platform also offers a social aspect; users have various ways of interacting with others on Spotify, from following the activity of their friends, sharing music, to collaboratively building playlists together. For this study, we use datasets describing the listening and sharing history of users on the platform, as well as their social activity with other users.
Figure 1: A synthetic network representing the structure of Spotify social network. The Spotify social network is a multiplex network, each layer corresponding to a different social interaction. Social listening sessions and collaborative playlists represent undirected layers, while link sharing is a directed layer.
@@
The Spotify social network comprises its users (nodes) and their interactions on the platform (edges). Each edge has an associated timestamp, indicating when an interaction occurred. This network is multiplex, each layer representing a different type of interaction; collaborative playlists, social listening sessions and link shares. A collaborative playlist is a special type of playlist that can be modified by a group of users. Although these are group interactions, we represent them as multiple pairwise interactions (edges) between each pair of collaborating users. The edges representing these interactions are undirected, as they are a reciprocal interaction. A social listening session allows multiple users to add songs to a real-time queue, listening to the same music either remotely or together. Like collaborative playlists, these are reciprocal interactions, and the edges are undirected. A link share is a link to a track, album or artist sent from one user to another via a messaging or social media platform. This interaction is directed, and is represented in the network by an edge from the sender to the receiver.
Figure 1 gives a depiction of this layered structure on a synthetic network. Each layer can be described by an adjacency matrix $L$, an $N\times N$ matrix where $N$ is the number of users. The values of the matrix $L_{ij}$ are the number of interactions occurring from user $i$ to user $j$. In the case of an undirected layer, where the interactions are reciprocal, the matrix is symmetric and $L_{ij}=L_{ji}$.
The resulting social network has three layers:
  • (1) Social listening sessions, an undirected layer with adjacency matrix $L_{s}$.
  • (2) Collaborative playlists, an undirected layer with adjacency matrix $L_{c}$.
  • (3) Link sharing, a directed layer with adjacency matrix $L_{l}$.
When aggregating the adjacency matrices of all the layers as $L_{s}+L{c}+L_{l}$, the network has $\sim$2.2B edges between $\sim$518M users (nodes). We specify different aggregations of these layers in the results sections, depending on the metric we are computing.
@@
We only consider “discovery” share events where the receiver has not listened to the artist before and we filter the data to include only link shares of tracks for which 1) the link was opened (comprising 50% of all shares), and 2) the receiver’s Spotify account registered a playback event of the shared content for at least 30 seconds (64% of all opened shares). This ensures that the receiver did indeed come into contact with the music that was shared. To account for popularity bias, we randomly sample up to 1M share events stratified by artist across various popularity bins. Our final sample consists of $\sim$2.72M events across 3,102 distinct artists between April-June 2023. Each link sharing event contains:
  • Anonymised sender and receiver identifiers.
  • Track and album identifiers, and the primary artist identifier.
  • Artist popularity rank at time of share.
  • Album release age from time of share.
  • Social media application type (e.g. WhatsApp).
  • Share timestamp and share open timestamp.
The identifiers and timestamps define the link sharing events, which are the basis of our analysis. The artist popularity, album age and application type are features in our analysis and predictive model, also described in Table 1.
We split the social media application type into two categories, direct or broadcast, depending on whether they represent one-to-one or one-to-many modes of communication. The applications we identify as direct are: WhatsApp, Facebook Messenger, SMS, Line, Instagram Direct, and Samsung Messenger. The applications we identify as broadcast are: Instagram Stories, Facebook Feed, Facebook Stories, and X, formerly Twitter. Some sharing events have an unknown destination platform or fit into neither category. These events are excluded from the analysis where share category is relevant. Note that the applications we identify as one-to-one can also be used to message group chats, but the key distinction here is in the magnitude of number of people receiving the link. Figure 2 gives the distribution of album release age at share-time for our link sharing event sample, separated by direct and broadcast integration. We can see that older tracks are more likely to be shared because there is more old content than new on the platform.
Figure 2: Distribution of the album release age at share-time for our link sharing events, separated by broadcast and direct application types.
@@
To this end, we use vector representations of Spotify users and content in latent space, which are pre-trained and available internally at Spotify. These vector embeddings are computed daily on Spotify by training the word2vec algorithm (Mikolov et al. 2013) on user generated playlists, with the playlists as the “documents” and the tracks as the “words”. This process results in 80-dimensional embedding vectors for each track, where tracks that are often co-listed in playlists appear close together in the embedding space. Users are also embedded in this space by taking the average of vectors over all tracks they have listened to over a certain time window. For more information on how these vector embeddings are generated, see (Anderson et al. 2020). The information captured by the embedding space methodology has been reported to be a good representation of the way that users engage with music (Mok et al. 2022; Anderson et al. 2020). We use vectors computed at the beginning of our analysis period for each user and track entity in our sample, to ensure consistency. The vectors can be used to understand the taste similarity of two users, or between a user and track. We define this as the cosine similarity between the two vectors.
@@
To define the success of a share, we are interested in a user’s engagement with an artist before and after a share event. We denote the number of unique tracks by the artist $\alpha$ that the user $i$ streams for longer than 30s on a particular day $t$ as $n_{i,\alpha}(t)$. We then define the user-artist engagement as $e_{i,\alpha}(t)=\log_{10}(n_{i,\alpha}(t))+1$, if $n_{i,\alpha}(t)\geq 1$, and 0 otherwise. User-artist engagement captures the intensity of the user’s engagement with the artist each day, which we aggregate over time as:
@@
where $t_{0}$ is the day when the link share is opened, and $K$ is the number of days over which we aggregate, which can be either positive or negative indicating aggregation over time before or after $t_{0}$ respectively. There are several choices that can be made to measure user-artist engagement. For example, the total number of streams over a window of time gives a broad indication of interest, the number of unique tracks listened to over a window of time gives an indication of deep interest, or the number of days listened to an artist, which is a longitudinal metric indicative of loyalty. Our choice of $E$ in Eq. (1) balances all three by looking at the unique number of daily tracks listened over a time window. The deeper (more daily tracks) and extended (more days) the listening is, the higher the engagement. The advantage of using this metric, instead of simply counting the number of unique tracks that the user listens to over the same period of time, is that $E_{i,\alpha}(t_{0},K)$ is higher if the number of listening events are dispatched over a larger number of days. For example, Fig. 3 (a) shows the profile of $E$ for a varying number of daily unique tracks listened $n_{i,\alpha}(t)$ and the total number of days listened $t_{\max}$ within a 7 day period. As we increase the number of days of engagement and the number of tracks listened, we define different curves for the engagement which reward high and repeat engagement. This behaviour is desirable when trying to capture the development of a habitual relationship such as becoming a fan of an artist. Thus, we capture the delineation between a state of longer-term engagement, and an intense, but time-limited, fixation with the artist. We use this metric to define the past engagement of a sender and the future engagement of a receiver with the shared artist.
Figure 3: User-artist engagement threshold choice. (a) User-artist engagement curves as a function of the number of tracks $n_{i,\alpha}(t)$ for different number of days listened $t_{\max}$ within a 7 day period. (b) Cumulative distribution function (ecdf) of the link shares and the associated receiver-artist engagement 7 days post share ($E_{i,\alpha}(t_{0},7)$. The orange line indicates our chosen threshold for a successful share event.
We define the success of a share event as the binary outcome of the receiver’s future engagement with the shared artist. Figure 3 (b) gives the cumulative distribution function of $E_{i,\alpha}(t_{0},7)$, the receiver-artist 7 day post-share engagement. We observe that $52.8\%$ of shares result in an engagement of 1.0, indicative of the user only listening to the shared track once in the 7 days post-share. Consequently, we define an engaged receiver as anyone who has listened to two or more tracks by the artist in the 7 days following the share, this is equivalent to a threshold of $E_{i,\alpha}(t_{0},7)>1.3$, since $\log_{10}{2+1}=1.3$. This threshold is shown in Fig. 3 (b) by the dotted orange line. Intuitively, exploring an artist’s other tracks and returning to the artist’s music after the initial share is a good sign that the user is engaging with the artist and the recommendation is successful, and so this choice of threshold identifies receiving users who are beginning to form an attachment with the artist in the short-term and are becoming potential fans. By this definition, $47.2\%$ of link shares result in successful engagement, and fairly balanced split between the link share events. Note that there are other choices that could be made to define a successful share. In particular, here we are interested in behaviour change on a short-term timescale directly following a share.
@@ Results
We evaluate the three hypotheses set out at the beginning of the paper by examining the probability of engagement $p(\mathrm{engaged})$, i.e. the proportion of link share events where the receiver’s engagement is $E_{i,\alpha}(t_{0},7)>1.3$ out of all share events, with respect to: the music taste similarity of the pair of sharing users and the shared track (H1), their relationship strength (H2), the social cohesion of their neighbourhood with respect to an artist (H3). We then construct and evaluate a predictive model of new listener engagement, using a broader set of social and contextual features.
@@
We measure music taste similarity as the cosine similarity of the music vectors described in Data and Methods. We first observe the level of homophily present in the sample of people sharing links, whether successful or not. We randomly shuffle the sender in user pairs in our sample, creating random permutation pairs of the social network, and take the user pair cosine similarity of their taste vectors. Figure 4 compares the distribution of the user pair cosine similarities between random pairs, and between the pairs who have shared and opened links, where we note that the distribution of user pair cosine similarities for users involved in link sharing is skewed towards higher similarity with a mean of 0.54, compared to the random pairs with a mean of 0.11. The difference in distributions between the two groups is statistically significant as confirmed by the Kolmogorov-Smirnov test ($D=0.44,p<0.01$). This gives some indication of the level of homophily in the Spotify social network; either users are more likely to become friends with other users who share similar tastes, or they are more likely to share with their friends who have similar tastes.
For this reason, it is difficult to determine whether the engagement we observe following a share is due to homophily or social contagion. However, while we may not be able to distinguish between the two mechanisms at play, we can still investigate the social and contextual factors that lead to engagement as a new listener of the shared artist. We first test the hypothesis (H1) that the probability of engagement will increase when users have similar taste in music (sender-receiver cosine similarity), or when the shared track is similar to the listening history of the receiver (receiver-track cosine similarity). The results are given in Fig. 5 (a), showing the distribution of receiver-track cosine similarity in our sample, and $p(\mathrm{engaged})$ as a function of the similarity. Similarly, Figure 5 (b) shows the results for sender-receiver cosine similarities. We observe that while (H1) is confirmed, there is a stronger effect for the receiver-track cosine similarity; $p(\mathrm{engaged})$ increases linearly with cosine similarity.
@@
We also investigate the effect of the sender-artist engagement for the artist on $p(\mathrm{engaged})$. We measure this by taking the 7 day aggregate user-artist engagement (described in Data and Methods) for the week prior to the share, which is a signal for how engaged the sender was with listening to the artist, prior to sharing their music. This corresponds to Equation 1, where $i$ is the sender, $t_{0}$ is the day of the share, and $K=-7$. This feature can be thought of as a proxy for the strength of the sender’s recommendation. Users have the ability to add messages when sharing a Spotify link, and while we do not have access to this textual data, it is possible that the associated message might be more emphatic when the sender is more engaged with the artist. Note that the sender-artist engagement can be 0 if the sender sends a track to someone without listening to it first, as these are separate functions on the Spotify interface. These results are shown in Fig. 6. The probability of engagement of the receiver increases with the sender-artist engagement.
@@
As suggested in (Almaatouq et al. 2016) and (Bond et al. 2012), stronger social ties are more effective at spreading behaviours. We explore the hypothesis (H2) that the probability of engagement $p(\mathrm{engaged})$ will increase with stronger and more intimate ties, as represented by the tie strength from the social network. We measure tie strength in three ways; reciprocity, the mode of social media application used to share the track, and sum of social interactions. Reciprocity, is a boolean variable that indicates whether the receiver ($i$) has ever shared a link with the sender ($j$), prior to share time. Specifically, we measure if $L_{l;ij}>0$. Figure 7 (a) shows the distribution of sender-receiver cosine similarities over all share events, separated by reciprocal and non-reciprocal ties. For reciprocal links, the mean sender-receiver cosine similarity is 0.58, compared to 0.5 for non-reciprocal links. The difference in distributions between the two groups is statistically significant as confirmed by the Kolmogorov-Smirnov test ($D=0.12,p<0.01$). This indicates that users connected by reciprocal ties tend to have a greater taste similarity than those who do not. Figure 7 (b) shows the probability of becoming an engaged receiver, $p(\mathrm{engaged})$, along a reciprocal and non-reciprocal link, holding constant the receiver-track cosine. Even accounting for the taste of the receiver, the probability of future engagement is more likely along a reciprocal tie. There is noise at very large and very small receiver-track cosine similarities, where there is little data available.
@@
Figure 7 (c) shows the distribution of sender-receiver cosine similarities over all share events, separated by ties that represent direct or broadcast platform share events. For the users sharing via direct social media applications, the mean user-user cosine similarity is 0.54, compared to the 0.57 for users sharing via broadcast methods. The difference in distributions between the two groups is statistically significant as confirmed by the Kolmogorov-Smirnov test ($D=0.03,p<0.01$). This indicates that users selectively engage with music shared by friends with similar taste on social media. Figure 7 (d) shows the probability of becoming engaged with the artist for share events via direct and broadcast mediums, again holding constant the user-track cosine similarity. Accounting for the taste of the receiver, the future engagement with an artist is more likely to occur when the mode of communication is a direct one, even when the track is not as good a match for the receiver’s taste. Note that the probability of engagement goes almost to 0 for broadcast application types at large receiver-track cosine similarities.
@@
The final measure of tie strength, sum of social interactions, the number of times that the sender ($j$) and receiver ($i$) have interacted before the share on Spotify. This includes link shares (from the sender to receiver or from the receiver to the sender), social listening sessions and blends. The sum of social interactions is equivalent to $L_{s;ij}+L_{c;ij}+L_{l;ij}+L_{l;ji}$, where we have included the link sharing in both directions. This measure gives a sense of whether the social tie is new, or previously established. Figure 8 shows the distribution of the sum of social interactions over the pairs of users in our sample, as well as the probability of engagement along a link as a function of the sum of social interactions. The probability of engagement jumps from 0.44 to 0.48 just after one previous interaction and then gradually up to 0.5, until 5 interactions. After 5 interactions, the probability of engagement remains stable. This indicates that while engagement is more likely to happen along an established tie, there is a “saturation point” at which more previous interaction does not increase the probability.
@@
Next, we explore the effects of social cohesion. These results investigate how the level of engagement with the shared artist in the receiver’s neighbourhood of contacts affects the probability of becoming engaged as a receiver of a share link (H3). We define the receiver’s “friends” as anyone with whom they have a reciprocal relationship in the social network. This includes social listening sessions, collaborative playlists, or link sharing (must be in both directions). More specifically, the receiver $(i)$ is friends with a user $(u)$ if $L_{s;iu}>0$, or $L_{c;iu}>0$, or $L_{l;iu}>0$ and $L_{l;ui}>0$. Figure 9 (a) shows the distribution of number of friends over receivers in the sample. The average number of friends is 3.95, but the distribution is skewed to the left, most users have 0 or 1 friend. A receiver’s friend is considered to be engaged with the shared artist if they have a user-artist engagement indicative of listening to at least one track per day on average by the shared artist in the six months prior to the analysis period. This is equivalent to the condition $E_{i,\alpha}(t_{0},180)>180$, where $i$ is the friend user, $\alpha$ is the shared artist and $t_{0}$ is the start of the analysis period. We expect that the probability of engagement of a receiving user will increase with the number of friends that the receiver has, who are already engaged with the artist, as a consequence of either social influence, or homophily in the social network.
We group all receivers who have the same number of friends, and track the probability of engagement as a function of the number of engaged friends of the receiver. The reason for holding constant the number of friends is to control for the absolute number of friends, as well as the fraction who are engaged. The results are shown in Fig. 9 (b) for 2,3 and 5 friends. The shaded area gives the $95\%$ confidence interval. In all cases, the probability of engagement increases with the number of engaged friends, supporting the hypothesis that social cohesion increases the probability of engagement.
It is possible that this effect that we observe is related to the artist popularity. In particular, the fraction of listener friends who are engaged with an artist may simply be a proxy for popularity, as these artists have more engagement overall. To disentangle these effects, we separate out the artist according to their popularity on Spotify, and plot the probability of engagement as a function of number of engaged friends. The results are shown in Fig. 9 (c). The probability of engagement increases with the number of engaged friends, even when accounting for artist popularity. More interestingly, we notice that this effect is more pronounced for the less popular artists, perhaps indicating that such social effects play a more important role in music discovery for lesser known artists.
@@
We would also like to evaluate the insights from H1-3 together with other contextual factors which might influence short-term engagement following a share. To this end, we design a supervised binary classification task, using data available at share-time to predict engagement in the week following the share. This allows us to elucidate further the social and contextual factors which explain whether a shared track will result in a successful person-to-person music recommendation.
@@
Sum of social interactions Sum of the number of social interactions between the sender ($j$) and receiver ($i$) prior to share-time. Includes collaborative playlist additions, social listening sessions, and link shares between the two users in either direction. Sum of social interactions = $L_{s;ij}+L_{c;ij}+L_{l;ij}+L_{l;ji}$.
Direct link share Boolean, True if the social media application type used to share the link was direct (e.g. Whatsapp), as opposed to broadcast or unknown.
Reciprocal link sharing Boolean, True if receiver has ever shared a link with the sender prior to share. This is equivalent to $L_{l;ij}>0$.
@@
Receiver link share in-degree Number of link shares received by the receiver ($i)$ (from anyone) prior to share-time. This is equivalent to $\sum_{k}L_{l;ki}$
Receiver link share out-degree Number of link shares sent by receiver ($i$) prior to share-time (to anyone). This is equivalent to $\sum_{k}L_{l;ik}$.
Sender link share out-degree Number of link shares sent by the sender ($j$) before the share-time (to anyone). This is equivalent to $\sum_{k}L_{l;jk}$.
@@
Fraction of engaged friends “Friend” is defined as someone with whom the receiver has a reciprocal relationship in the social network prior to share-time. A friend is considered engaged if $E_{u,\alpha}(t_{0},180)>180$, where $u$ is the friend user, $\alpha$ is the shared artist and $t_{0}$ is the start of the analysis period.
@@
sender-receiver cosine similarity Cosine similarity between receiver and sender.
receiver-track cosine similarity Cosine similarity between receiver and track.
@@
Artist popularity at share-time Rank of the primary artist in Spotify by number of streams at the time of share.
Release age at share-time Time elapsed between album release time and share-time (seconds).
@@
Sender-artist engagement 7 days prior Sender-artist engagement aggregated over 7 days prior to share-time.
@@
is subscriber Whether the user was a paid subscriber to Spotify when the share was opened.
receiver streaming 7 days prior Number of hours of streamed music in the previous 7 days to opening the share.
receiver day on platform Number of days passed since the user registered to the platform to the day the share was opened.